2x+5y = -1
x+2y = 0

Found 2 solutions by gsmani_iyer, kiru_khandelwal:
Answer by gsmani_iyer(201)


2x + 5y = -1 ..... (1)
x + 2y = 0 ..... (2)
(2) * 2, we get (To make one term in both the equations
2x + 4y = 0 .....(3) equal.)
(1) - (3), we get
y = -1 .........(4)
Substituting the value of y in Eq.(1), we get
2x + (-5) = -1
So 2x = 5 - 1 = 4
So x = 4/2
= x = 2.
So x=2; y=-1; Answer.

Answer by kiru_khandelwal(79)

x+2y = 0
=> x = -2y............Equation (1)
2x + 5y = -1
Substituting the value of x from equation (1)
2(-2y) + 5y = -1
=> -4y + 5y = -1
=> y = -1
Substituting the value of y in equation (1)
x = -2y = -2*(-1) = 2
x = 2 and y = -1